OpenStudy (anonymous):

Miss Moneypenny inherits £1050 from a secret agent she decides to save some of the money and spend the rest the ratio of savings to spending money is 7:8 how much does she save and how much does she spend Savings..... Spending Money....

OpenStudy (anonymous):

sav = 1050 - spe sav/spe = 7/8 (1050 - spe)/spe = 7/8 spe = 560 sav = 490

OpenStudy (anonymous):

A little bit more detail: 8(1050 - spe) = 7spe 8400 - 8spe = 7spe 8400 = 15spe spe = 560

OpenStudy (anonymous):

Arthur is organizing his collection of DVD'S he has 120 altogether The Ratio of comedy films to action to horror is 2:8:5 How many of each type of DVD'S Does he have Comedy.....DVD'S Action......DVD'S Horror.......DVD'S

OpenStudy (anonymous):

Could you help me with that one

OpenStudy (anonymous):

Is this making sense now or do you still have questions about this problem?

OpenStudy (anonymous):

@James32841234 ?

OpenStudy (anonymous):

Plz

OpenStudy (anonymous):

So you don't answer questions when a tutor asks you something?

OpenStudy (anonymous):

Plz Help me

OpenStudy (anonymous):

Arthur is organizing his collection of DVD'S he has 120 altogether The Ratio of comedy films to action to horror is 2:8:5 How many of each type of DVD'S Does he have Comedy.....DVD'S Action......DVD'S Horror.......DVD'S